Circular No. 7743

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ff IN UGC 4685 W. D. Li, University of California at Berkeley, reports the discovery

by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 17.1) in unfiltered KAIT images taken on Nov. 2.5 and 3.5 UT. SN 2001ff is located at α = 8^{h}57^{m}00^{s}.97, δ = +13^{o}11′59′′.5 (equinox 2000.0), which is 1′′.5 west and 2′′.6 north of the nucleus of UGC 4685. A KAIT image taken on Apr. 17.2 showed nothing at this position (limiting mag ∼ 19.0).

NOVA IN THE SMALL MAGELLANIC CLOUD 2001 M. Della Valle, Arcetri Astrophysical Observatory, Florence; I. Sa-

viane, European Southern Observatory; and R. Williams, Space Telescope Science Institute, report that preliminary analysis of a spectrum (range 390–850 nm), obtained on Oct. 30.25 UT at La Silla with the 1.5-m telescope (+ Boller & Chivens spectrograph), shows that the suspected nova reported on IAUC 7738 is indeed a nova caught during its early decline. The spectrum is dominated by emission lines of the Balmer series, flanked by P-Cyg profiles. Also prominent are emission lines due to Na I, Fe II, and O I. The FWZI measured for Hα yields ∼ 5500 km/s.

E. L. N. Jensen, Swarthmore College; P. Allen, University of Pennsylvania; G. J. Schwarz, University of Arizona; and K. M. Vanlandingham, Biosphere 2 Center, report on optical spectra (range 480–800 nm, resolution 0.014 nm) taken of the nova on Nov. 2.12 UT with the Cerro Tololo 4-m reflector (+ échelle spectrograph). The spectrum is dominated by strong emission lines of Hα, Hβ, and O I (777.1 nm). The Balmer lines have average FWHM = 1400 km/s and FWZI = 2200 km/s. Weak lines of C II (723.4 nm), [O I] (630, 636.3, and 557.7 nm), and Fe II (multiplets 42, 74) are also seen. The lines all show a double-peaked structure. The spectrum is very similar to that of N LMC 1988 No. 1, ≈ 2 weeks after outburst.

SUPERNOVA 2001fc IN UGC 11683 M. Ganeshalingam and W. D. Li, University of California at Berkeley,

report the independent discovery of SN 2001fc (at mag ∼ 18.7) by LOTOSS on an unfiltered KAIT image taken on Oct. 27.1 UT. They measured the position end figures for SN 2001fc (cf. IAUC 7741) to be 26^{s}.66, 15′′.6. A KAIT image taken on Oct. 13.1 showed nothing at this position (limiting mag ∼ 19.0).

2001 November 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green